I'm not too sure if this is the right section. But here goes.
Simplify the following:
(n^5/m^2)^3 x (m^4/n^2)^4

Answer by uday1435(57)

(n^5/m^2)^3 x (m^4/n^2)^4
= (n^15/ m^6) x (m^16/n^8)
= n^(15-8) m^(16-6) = n^7 m^10
So the answer is n^7 m^10


If you still require further clarifications please write to udayakumar.t.r@gmail.com